HIGH RESOLUTION INFRARED ABSORPTION SPECTROSCOPY OF THE FIRST OVERTONE PURE VIBRATIONAL TRANSITION Q_{2}(0) OF SOLID PARAHYDROGEN

\underline{HIROYUKI KATSUKI} and TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, JAPAN.

It has been shown that high-resolution infrared spectroscopy is applicable to the zero-phonon transitions of rotons and vibrons in solid hydrogen. Because the pure vibrational transitions Q_{n}(0) [v=n←0, J=0←0] of solid parahydrogen become infrared-active only upon interaction with residual J=1 orthohydrogen, studies of the pure vibrational transitions under high-resolution provide detailed and accurate information on intermolecular interactions between hydrogen molecules. Here, we report the high-resolution absorption spectrum of the first overtone pure vibrational transition Q_{2}(0) of solid parahydrogen. The high-resolution spectrum of the Q_{2}(0) transition at around 8070 cm^{−1} was observed using a difference frequency laser system^{a}. The spectrum shows a complicated spectral feature compared with that of the Q_{3}(0) transition^{b}. The v=3 vibrational exciton was found to be well localized on the parahydrogen molecule next to an orthohydrogen. On the other hand, since the vibron hopping matrix element of the v=2 state is comparable to the energy shift due to the existence of orthohydrogen, the observed rich spectral structure can be interpreted as due to the v=2 vibron hopping over a limited number of lattice sites in the crystal. We will discuss the quantitative analysis of the observed spectrum.

aT. Momose, T. Wakabayashi, and T. Shida, J. Opt. Soc. Am. B 13, 1706 (1996).

bR. M. Dickson, T. Momose, T. J. Byers, and T. Oka, Phys. Rev. B 57, 941 (1998).